# HSF: Defending against Jailbreak Attacks with Hidden State Filtering

### v1

@@ 5.1 Experimental Setup
5.2 Experimental Results
HSF Enhances LLM Safety. Table 2 presents a comparison of the ASR for the Mistral-instruct-v0.2 and Llama2-7B-chat models when employing HSF and various baseline defenses against six different jailbreak attacks. The results indicate that for models with weaker safety alignment, such as Mistral-instruct-v0.2, HSF markedly decreases the ASR, outperforming almost all baseline defenses. Notably, while most other defenses were ineffective against AutoDAN, HSF successfully defends against this attack, achieving an ASR of 0. For models with strong alignment, such as Llama2-7B-chat, HSF consistently reduces the ASR across all attacks to nearly 0. Additional results for HSF applied to Vicuna-v1.5 can be found in AppendixB.2.
@@ 5.2 Experimental Results
Table 3: Comparison of AUC for Mistral-instruct-v0.2 and Llama2-7B-chat across different K values (1-8).
5.3 Ablation Analysis
@@ 6 Conclusion
7 Limitations
@@ Appendix A Appendix A: Detailed Experimental Setups
A.1 Attack Setup
For GCG (Zou et al. 2023), AutoDAN (Liu et al. 2024), GPTFuzz (Yu, Lin, and Xing 2023), DeepInception (Li et al. 2023), and ICA (Wei, Wang, and Wang 2023), we use the easyjailbreak (Zhou et al. 2024) framework to construct attacks. We utilized 100 different representative harmful queries from Advbench (Zou et al. 2023). For ReNellm (Ding et al. 2024), we rewrite these 100 representative harmful queries collected from Advbench using easyjailbreak and filled them into templates to create 100 jailbreak prompts.
A.2 Baseline Setup
1. PPL (Alon and Kamfonas 2023): PPL is an input detection mechanism that calculates the perplexity of a given input to determine whether the user’s request should be accepted or rejected.
@@ A.2 Baseline Setup
$\text{Perplexity}(P)=\exp\left(\frac{-1}{N}\sum_{i=1}^{N}\log P(w_{i}\mid w_{1:i-1})\right).$ (4)
Following (Alon and Kamfonas 2023), we use Llama2-7b-chat to calculate perplexity. As per (Xu et al. 2024), we set the PPL threshold to the highest perplexity value of harmful queries in Advbench. This ensures that queries from Advbench do not trigger the detector.
2. Paraphrase (Jain et al. 2023): We follow (Jain et al. 2023) and used GPT-4o by default to paraphrase user queries. The prompt is as follows:
@@
3. Retokenization (Jain et al. 2023): This method splits tokens and represents them using multiple smaller tokens. For example, the token representing the word “breaking” is split into two tokens representing “break” and “ing.” We use BPE-dropout (Provilkov, Emelianenko, and Voita 2020) , which randomly drops pof BPE merges during the tokenization process. We set p = 0.2 following (Jain et al. 2023).
@@
4. Self-Examination (Phute et al. 2023): This is an output detection method that uses the language model itself to distinguish whether harmful content is being generated. We use the prompt suggested by (Phute et al. 2023):
@@
5. Self-Reminder(Xie et al. 2023): Self-Reminder appends prompts to input prompts to remind the language model to respond responsibly.
@@
6. DRO (Zheng et al. 2024): This is an automated optimization method for generating safe prompts. Following previous work, we use a default type of system safety prompt and directly embedded the generated safety prompt as a vector in the embedding layer before the user’s input.
Appendix B Appendix B: More Results
B.1 PCA visualization in More Models
We also conducted PCA visualizations on two additional models in Figure4 and Figure5.
Figure 4: Comparison of Mistral-7B-Instruct-v0.2 model using PCA visualization, illustrating how the model differentiates between harmless (blue) and harmful (orange) queries across various jailbreak methods. Subplots correspond to the following attack methods: (a) Autodan, (b) GCG, (c) DeepInception, (d) ReNellm, (e) ICA, and (f) GPTFuzz. The decision boundaries that the model learned to separate the two types of queries are shown.
Figure 5: Comparison of Vicuna-7b-v1.5 model using PCA visualization, illustrating how the model differentiates between harmless (blue) and harmful (orange) queries across various jailbreak methods. Subplots correspond to the following attack methods: (a) Autodan, (b) GCG, (c) DeepInception, (d) ReNellm, (e) ICA, and (f) GPTFuzz. The decision boundaries that the model learned to separate the two types of queries are shown.
B.2 Hidden State Filter in More Models
We demonstrate HSF when applied in Vicuna-v1.5 in Table 5. Our observations reveal that, although jailbreak attacks on Vicuna-v1.5 yield high ASR and harmful scores, HSF can significantly mitigate its effectiveness.
Vicuna No Defense 60% 51% 94% 100% 84% 77% 29% 51%
HSF 1% 5% 10% 0% 7% 0% 0% 0%
Table 5: HSF appied in Vicuna-v1.5.We note that HSF significantly mitigates the effectiveness of current state-of-the-art attacks in Vicuna-v1.5.
B.3 Jailbreak attack templates is essential
Incorporating adversarial attack templates into the training set has proven to be a highly effective strategy for enhancing the robustness of lightweight classification models against adversarial inputs (Wallace et al. 2019). By systematically exposing the model to a diverse array of adversarial prompts during training, we empower the model to more effectively identify and counteract such attacks during inference. This proactive exposure helps establish more resilient decision boundaries within the model’s representational space, thereby significantly reducing the success rate of jailbreak attempts.
To demonstrate the effectiveness of this approach, we present a comparative analysis of models trained with and without the inclusion of jailbreak attack templates in Table 6 . The results clearly show that models trained with these templates achieve a substantially lower Attack Success Rate (ASR) across various jailbreak methods, highlighting the critical importance of this training enhancement in improving model robustness.
Model GCG AutoDAN ReNellm DeepInception ICA GPTFuzz
Vicuna (w/o Jailbreak Templates) 16% 0% 68% 11% 0% 0%
Vicuna (w/ Jailbreak Templates) 10% 0% 7% 0% 0% 0%
Mistral (w/o Jailbreak Templates) 0% 0% 17% 93% 0% 0%
Mistral (w/ Jailbreak Templates) 4% 0% 12% 0% 0% 0%
Table 6: A comparative analysis of ASR for the Vicuna-v1.5 and Mistral-instruct-v0.2 models trained with and without jailbreak attack templates. The results demonstrate that models trained with these templates consistently exhibit lower ASR, indicating enhanced robustness against adversarial attacks.
This table underscores the substantial reduction in ASR when jailbreak templates are included in the training set, thereby highlighting the importance of such templates in building more secure and reliable LLMs.
Appendix C Appendix C: Example Demonstrations
We present the following examples illustrating HSF across different models. For clarity, attack prompts are highlighted in red.
This example shows HSF is applied in Llama2-7b-chat to defend against AutoDAN(Liu et al. 2024)
This example shows HSF is applied in Mistral-instruct-v0.2 to defend against Deepinception(Li et al. 2023)
This example shows HSF is applied in Vicuna-v1.5 to defend against ReNellm(Ding et al. 2024)

### v2

@@ 5.1. Experimental Setup
5.2. Experimental Results
HSF Enhances LLM Safety. Table 2 presents a comparison of the ASR for the Mistral-instruct-v0.2 and Llama2-7B-chat models when employing HSF and various baseline defenses against six different jailbreak attacks. The results indicate that for models with weaker safety alignment, such as Mistral-instruct-v0.2, HSF markedly decreases the ASR, outperforming almost all baseline defenses. Notably, while most other defenses were ineffective against AutoDAN, HSF successfully defends against this attack, achieving an ASR of 0. For models with strong alignment, such as Llama2-7B-chat, HSF consistently reduces the ASR across all attacks to nearly 0. Additional results for HSF applied to Vicuna-v1.5 can be found in AppendixLABEL:appendix:more_results.
@@ 5.2. Experimental Results
Table 3. Comparison of AUC for Mistral-instruct-v0.2 and Llama2-7B-chat across different K values (1-8).
@@
5.3. Ablation Analysis
Table 4. Ablation analysis on the hyperparameter $k$ for HSF.
@@ 7. Conclusion
Acknowledgments
This work was funded by the National Natural Science Foundation of China (NSFC) under Grants No. 62406013, the Beijing Advanced Innovation Center Funds for Future Blockchain and Privacy Computing and the Fundamental Research Funds for the Central Universities.
@@ Appendix A Appendix A: Detailed Experimental Setups
A.1. Attack Setup
For GCG (Zou et al., 2023), AutoDAN (Liu et al., 2024), GPTFuzz (Yu et al., 2023), DeepInception (Li et al., 2023), and ICA (Wei et al., 2023), we use the easyjailbreak (Zhou et al., 2024) framework to construct attacks. We utilized 100 different representative harmful queries from Advbench (Zou et al., 2023). For ReNellm (Ding et al., 2024), we rewrite these 100 representative harmful queries collected from Advbench using easyjailbreak and filled them into templates to create 100 jailbreak prompts.
A.2. Baseline Setup
1. PPL (Alon and Kamfonas, 2023): PPL is an input detection mechanism that calculates the perplexity of a given input to determine whether the user’s request should be accepted or rejected.
@@ A.2. Baseline Setup
(4) $\text{Perplexity}(P)=\exp\left(\frac{-1}{N}\sum_{i=1}^{N}\log P(w_{i}\mid w_{1:i-1})\right).$
Following (Alon and Kamfonas, 2023), we use Llama2-7b-chat to calculate perplexity. As per (Xu et al., 2024), we set the PPL threshold to the highest perplexity value of harmful queries in Advbench. This ensures that queries from Advbench do not trigger the detector.
2. Paraphrase (Jain et al., 2023): We follow (Jain et al., 2023) and used GPT-4o by default to paraphrase user queries. The prompt is as follows:
@@
3. Retokenization (Jain et al., 2023): This method splits tokens and represents them using multiple smaller tokens. For example, the token representing the word “breaking” is split into two tokens representing “break” and “ing.” We use BPE-dropout (Provilkov et al., 2020) , which randomly drops pof BPE merges during the tokenization process. We set p = 0.2 following (Jain et al., 2023).
@@
4. Self-Examination (Phute et al., 2023): This is an output detection method that uses the language model itself to distinguish whether harmful content is being generated. We use the prompt suggested by (Phute et al., 2023):
@@
5. Self-Reminder(Xie et al., 2023): Self-Reminder appends prompts to input prompts to remind the language model to respond responsibly.
@@
6. DRO (Zheng et al., 2024): This is an automated optimization method for generating safe prompts. Following previous work, we use a default type of system safety prompt and directly embedded the generated safety prompt as a vector in the embedding layer before the user’s input.
7. SPD (Candogan et al., [n. d.]): This is a method to detect adversarial samples by extracting a feature matrix $H$ from the logit outputs of a LLM. $H$ captures unnatural token distribution changes caused by attacks, using the top-$k$ logits for $r$ token positions.The authors originally used a Support Vector Machine (SVM) with an RBF kernel trained on $H$ for classification. However, during our reproduction, we encountered issues with the SVM and instead replaced it with a MLP. The hyperparameters include $r$ (number of token positions, set to 5), $k$ (number of top logits, set to 50).